# Retain or Reframe? A Computational Framework for the Analysis of Framing in News Articles and Reader Comments

### v1

@@ 6.1 Frame Classification
We train frame classifiers that predict sentence-level media frames for both comments and articles, following the framing schema defined in Section 5.1. To do so, we require suitable training data from both genres, and a robust evaluation setup.
Frame Classification Data
We source articles from the Media Frames Corpus (Card et al., 2015) and forum comments from Hartmann et al. (2019), splitting it by sentences and deriving sentence-level annotations from the original span-level annotations. To do so, we first map the original frame labels to our 9 classes; then we identify spans with matching labels from at least two annotators with a minimum three-word overlap. Following previous methods (Naderi and Hirst, 2017; Hartmann et al., 2019; Park et al., 2022), agreed spans were mapped to the sentence-level: sentences fully covered by agreed spans inherited corresponding frame labels. This results in approximately 63,000 labeled sentences across both genres, but predominantly from the MFC (only 625 sentences from the much smaller Hartmann et al. (2019) dataset). Further details on frame and topic distributions of the training data are provided in Appendix D, Table 7.
@@
To automatically classify frames in our news articles and comments, we fine-tuned three models — RoBERTa-Large (Liu et al., 2019b) (full fine-tuning), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), and Qwen2.5-7B-Instruct (Qwen et al., 2025) (both parameter-efficient fine-tuning) — using the sentence-level annotated data described above on a random 80/10/10 (training/dev/test) split, stratified by both frame and topic. Fine-tuning parameters are reported in Appendix A, and prompting details in Appendix B.
@@
In-Domain 0.66 0.65 0.64
Table 2: Macro-F1 scores of three fine-tuned LMs on MFC + Hartmann data.
@@
RoBERTa 0.59 0.77 0.83
Table 3: Macro-F1 scores of Lama cross-domain (X-Domain) and on FrAC articles (A) and comments (C).
@@
Table 2 shows little difference in the models’ performance on the 10% MFC/Hartmann test split described above, with RoBERTa-Large slightly outperforming the larger Qwen2.5-Instruct and LLaMA3.1-8B-Instruct. This result aligns with recent literature on frame classification, where BERT-based pre-trained models have consistently demonstrated good performance (Sajwani et al., 2024; Daffara et al., 2025). Since RoBERTa also requires substantially less fine-tuning time (approximately 20 minutes vs. 13 hours for PEFT-tuning of LlaMA and Qwen), all remaining experiments in this paper are based on the fine-tuned RoBERTa model.
To evaluate model generalizability to unseen topics, we train RoBERTa on all but one MFC topics, evaluating on the held-out topic, averaging over all five combinations. Table 3 (X-Domain) revealed an expected, albeit relatively small drop in RoBERTa performance (from 0.66 In-Domain in Table 2 to 0.59), suggesting a reasonable ability to generalize to unseen topics, which is important because the model will be applied to topics not present in the MFC training set.
Finally, we evaluated RoBERTa on the annotated FrAC corpus (Table 3, FrAC), i.e., a data set from different outlets, time points and distinct yet overlapping set of topics (Section 5 and Table 4). The model achieved strong performance on FrAC with an F1 score of 0.77 on articles and 0.83 on comments, demonstrating effective generalization to reader comments despite being trained primarily on formal news articles. This result validates our central methodological contribution: that sentence-level frame classification can successfully bridge the gap between formal news articles and informal reader comments.
Having validated RoBERTa’s performance on our gold-standard FrAC corpus, we then applied it to the full sentence-tokenized SOCC and NYT data sets (Section 4.1) for our large-scale framing effects analysis. [^9] We used the NLTK library https://www.nltk.org/ for sentence tokenization of article and comment texts.
@@ 6.2 Dominant Frame Detection
We analyze framing effects between articles and comments in terms of their respective dominant frames. We define the dominant frame as the most frequent sentence-level frame label in a given document. To reduce noise from near ties, we only assign a dominant frame if the most frequent frame meets both of the following criteria: it appears in at least 3 sentences and it represents at least 40% of all sentences in the document. We make an exception for single-sentence texts (comments), whose frame is automatically considered to be dominant. Articles and comments whose most frequent frame fails to meet these combined thresholds were excluded from further analysis.
@@ 6.3 Topic Extraction
We applied BERTopic (Grootendorst, 2022) separately to SOCC and NYT to extract topics from news articles, assigning each article a probability distribution across all identified topics. [^10] We report the full list of parameters used in Appendix C. We analyzed the initial hierarchical topic structures produced by BERTopic and manually merged closely positioned topics with small inter-topic distances and high thematic similarity. The refined topic model was then applied to reader comments. We ultimately assigned each article and each comment the single most likely topic predicted by BERTopic.
This process was conducted independently for both corpora, to avoid any bias in subsequent analyses. The extracted topics showed substantial thematic overlap across data sets, enabling cross-dataset comparison. We manually identified eleven overarching, semantically cohesive topics which we use in our downstream analysis (Table 4).
@@
To assess the quality of topic assignments, we randomly sampled 5 articles and 5 comments from each of the 11 topic groups, yielding a total of 110 texts. Two authors of the paper independently reviewed each assigned topic and judged whether it was acceptable or not. We observed a high degree of inter-annotator agreement (96%). The overall accuracy of the model’s predicted topics, calculated against the labels assigned by the annotators was found to be 93% (on both comments and articles).
@@ 6.4 Resulting Data Set
With our framework and validated implementation at hand, we next apply it to analyze framing effects across the SOCC and NYT data sets (Section 4.1) and eleven selected topics (Section 6.3). We only include article-comment pairs where both share the same predicted topic (excluding off-topic comments to study framing effects when discussions remain on-topic) and where dominant frames could be extracted for both articles and comments. We also filter out comments shorter than 5 words, and repeated texts. This final data set consists of nearly 3,000 articles and over 380,000 associated comments, as detailed in Table 4.
@@ 7 Framing Effects across Topics and Outlets
Using our corpus of articles and comment frames from two outlets, annotated with both dominant frames and labels, we next address our three research questions (Section 3) in turn.
Figure 2: News article frames are retained in reader comments to varying degrees (X- and Y-axes). The degree of retention correlates strongly across two media outlets: New York Times (NYT) and The Globe and Mail (SOCC).
@@
7.1 To What Extent are Frames Retained vs Changed?
We quantitatively investigated the dominant frames employed in both the news articles and corresponding reader comments across the New York Times (NYT; USA) and The Globe and Mail (SOCC; Canada), analyzing the extent to which the dominant frame of the article was retained in the associated reader comments.
@@ 7.1 To What Extent are Frames Retained vs Changed?
Overall, we find that NYT readers demonstrate stronger frame retention (50.7%) compared to SOCC readers (37.3%). Conversely, at least half of the comments (almost two thirds in SOCC) change away from the dominant article frame indicating that readers often do not retain the original article framing, but change the message. In section 7.3, we analyze these reframing patterns in depth.
Figure 2 displays the frame retention patterns separately for each dominant article frame. Looking at the spread of frames along the x- and y-axis independently, we can see that the extent of retention varies by frame. Specifically, frames such as "Political and Policies", "Economic", and "Health and Safety" show high retention rates. In contrast, "Morality", "Fairness and Equality", and "Cultural Identity" exhibit substantially lower retention rates, suggesting that readers are more likely to reframe content that is presented through frames that require alignment with personal beliefs or values to be accepted (Chong and Druckman, 2007).
[…]
Figure 2 additionally reveals that the reframing patterns are remarkably consistent across NYT and SOCC, i.e., all points cluster close to the diagonal line of perfect correlation. This suggests that our uncovered patterns might reveal a general trend in news reader behavior, rather than being idiosyncratic to a particular outlet or sample, although future analyses across more diverse outlets and countries will need to verify this trend.
@@
Figure 3: Frame retention rates by topic across news outlets. Points above the diagonal line indicate higher retention in NYT, while points below indicate higher retention in SOCC. Topic names are shortened for enhanced visibility.
7.2 Do Framing Effects Vary Across Topics?
Here, we explore the effect of topics on framing effects. We cannot estimate topic effects from the observed data directly, because the effect of topics is confounded by the topic-specific frame distributions. To address this, we estimate the effect of article topic on frame retention using logistic regression mixed effects models. Given a document-response pair $i=(d,r)$:
$y_{i}\sim\text{topic}_{d}+(1|\text{frame}_{d})+(1|\text{id}_{d}),$ (1)
where $y_{i}$ is a binary indicator of frame retention ($\text{frame}_{d}{==}\text{frame}_{r}$). Document (article) topic ($\text{topic}_{d}$) is a fixed effect of topic on frame retention. And $(1|\text{frame}_{d})$ and $(1|\text{id}_{d})$ are fixed effects, controlling for the impact of the article frame ($\text{frame}_{d}$) and article id $(\text{id}_{d})$, respectively. [^11] We control for article ID due to repeated measurements, given that we have multiple comments per article. We then obtain the marginal effects of $\text{topic}_{d}$ on $y$ from the fitted model, removing any influence of the random effects. The resulting probabilities of frame retention per topic are shown in Figure 3 for NYT (y-axis) and SOCC (x-axis).
We find that framing effects vary by topic, but the degree of variance is much higher for NYT compared to SOCC suggesting that these patterns are more outlet specific than the framing-based variation discussed above and shown in Figure 2. This offers an explanation for the inconsistent findings in previous single-topic studies. Some topics exhibit consistent retention patterns across outlets: Coverage of Trump & Elections demonstrates relatively high retention in both data sets (NYT: 31.3%, SOCC: 42.7%), while Healthcare and Education also show moderate to strong retention. International conflict topics are consistently more often reframed by the readers: Israel-Palestine conflict and Syria & IS show the lowest retention rates across both sources, ranging between 18-24%. Other topics reveal outlet-specific differences: Immigration shows substantially higher frame retention in NYT than SOCC, as well as the coverage of the Russia-Ukraine conflict.
7.3 How Do Readers Reframe?
Comments change away from the dominant article frame 50% (NYT) or 63% (SOCC) of the time, so a natural question is how readers reframe articles. Figure 4 summarizes the flow of reframing in the NYT across all topics. Readers predominantly reframe issues into political terms, regardless of the original article framing, but a particularly large proportion of value-laden article frames (“Fairness and Equality”, “Public Opinion”) are recast as “Political and Policies”. [^12] This does not necessarily mean, though, that readers “politicize” all articles – “Political and Policies” frame is also the most reframed one, and its relative proportion in comments is actually lower than in the articles. This pattern, albeit to a lesser degree, also emerges in SOCC (Figure 6 in Appendix), where readers also often recast different frames into “Economy”.
We also investigated topic-specific patterns of reframing for topics that included at least 100 articles. For instance, in the case of Immigration (Table 5), SOCC readers both reframe “Cultural Identity” into “Political and Policies” terms, and vice versa, suggesting a tight interconnection of these two perspectives. NYT readers, on the other hand, primarily reframe into ‘harder’ perspectives like “Economic” and “Legality and Crime” and interestingly, among these top five reframings (as shown in the table), the “Economic” framing occurs only on the reader side. Table 9 in the Appendix shows the analysis for three additional topics, revealing varying trends across topics and outlets.
Figure 4: Reframing patterns in the New York Times showing the flow from primary article frames (left) to primary comment frames (right) with proportion of frames in articles and comments, respectively. We plot standardized frame proportions in articles, to reduce visual clutter.
Reframing Pattern N
SOCC Cultural Identity → Political & Policies 134
Political & Policies → Cultural Identity 131
Health & Safety → Cultural Identity 120
Economic → Political & Policies 108
Health & Safety → Political & Policies 99
NYT Legality & Crime → Political & Policies 1388
Political & Policies → Economic 1311
Political & Policies → Legality & Crime 1139
Legality & Crime → Economic 1024
Political & Policies → Cultural Identity 950

### v2

@@ 6.1 Frame Classification
We train frame classifiers that predict sentence-level media frames for both comments and articles, following the framing schema defined in Section 5.1. To do so, we require suitable training data from both document types, and a robust evaluation setup.
Training Data
As our FrAC corpus is too small to support fine-tuning, we derive large-scale labeled data sets from the MFC (news articles) and Hartmann et al. (2019) (comments). We sentence-split both datasets and derive sentence-level annotations from the original span-level annotations: we first map the original frame labels to our 9 classes; then we identify spans with matching labels from at least two annotators with a minimum three-word overlap. Following previous methods (Naderi and Hirst, 2017; Hartmann et al., 2019; Park et al., 2022), agreed spans were mapped to the sentence-level: sentences fully covered by agreed spans inherited the span labels. This results in approximately 63,000 labeled sentences predominantly from the MFC (including 625 sentences from the much smaller Hartmann et al. (2019) dataset). Further details on frame and topic distributions of the training data are provided in Appendix D, Table 9.
@@
To automatically classify frames in news articles and comments, we fine-tuned three models – RoBERTa-Large (Liu et al., 2019b) (full fine-tuning), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), and Qwen2.5-7B-Instruct (Qwen et al., 2025) (both parameter-efficient fine-tuning) – using the sentence-level annotated data described above with a random 80/10/10 (training/dev/test) split, stratified by frame and topic. Fine-tuning parameters are reported in Appendix A, and prompting details in Appendix B.
@@
P R F1 P R F1 P R F1
0.70 0.65 0.66 0.68 0.65 0.65 0.64 0.64 0.64
Table 3: Precision (P), recall (R) and Macro-F1 scores for the three fine-tuned LMs on MFC + Hartmann data.
Label P R F1
Economic 0.77 0.83 0.80
Morality 0.70 0.70 0.70
Fairness and Equality 0.68 0.59 0.64
Legality and Crime 0.83 0.86 0.85
Political and Policy 0.77 0.85 0.81
Security and Defense 0.69 0.59 0.63
Health and Safety 0.73 0.80 0.76
Cultural Identity 0.71 0.54 0.61
Public Opinion 0.61 0.61 0.61
Other 0.53 0.14 0.22
Table 4: Per-frame precision, recall, and Macro-F1 scores for fine-tuned RoBERTa on MFC + Hartmann data.
@@
P R F1 P R F1 P R F1
0.65 0.59 0.59 0.74 0.81 0.77 0.84 0.85 0.83
Table 5: Precision, Recall and Macro-F1 scores of RoBERTa cross-domain (X-Domain) and on FrAC articles (A) and comments (C).
@@
The averaged results in Table 3 show little performance difference across models, with RoBERTa-Large slightly outperforming the larger Qwen2.5-Instruct and LLaMA3.1-8B-Instruct. This result aligns with recent literature on frame classification, where BERT-based pre-trained models have consistently demonstrated strong performance (Sajwani et al., 2024; Daffara et al., 2025). Since RoBERTa also requires substantially less fine-tuning time (approximately 20 minutes versus 13 hours for PEFT-tuning of LlaMA and Qwen), all remaining experiments in this paper are based on fine-tuned RoBERTa. Table 4 reports the per-label breakdown for the best performing model, RoBERTa.
To evaluate model generalizability to unseen topics, we train RoBERTa on all but one MFC topics, evaluate on the held-out topic, average over all five combinations. Table 5 (X-Domain) revealed an expected, albeit relatively small drop in RoBERTa performance (from 0.66 In-Domain in Table˜4 to 0.59), suggesting a reasonable ability to generalize to unseen topics, which is important because the model will be applied to topics not present in the MFC training set.
Finally, we evaluated RoBERTa on the annotated FrAC corpus (Table 5, FrAC), i.e., a data set that differs from the MFC fine-tuning data in its outlets, time points and distinct yet overlapping set of topics (Section˜5 and Table˜6). The model achieved strong performance on FrAC with an F1 score of 0.77 on articles and 0.83 on comments, demonstrating effective generalization to reader comments despite being trained primarily on formal news articles. [^9] FrAC was annotated directly at the sentence level using the consolidated 9+1 frame taxonomy, whereas MFC labels were created post-hoc from span-level annotations, introducing label noise which likely negatively affected model performance. This result validates our central methodological contribution: that sentence-level frame classification can successfully bridge the gap between formal news articles and informal reader comments.
Having validated RoBERTa’s performance on our gold-standard FrAC corpus, we then applied it to the full sentence-tokenized SOCC and NYT data sets (Section˜4.1) for our large-scale framing effects analysis. [^10] We used the NLTK library (https://www.nltk.org/) for sentence tokenization of article and comment texts.
@@ 6.2 Dominant Frame Detection
We analyze framing effects between articles and comments in terms of their respective dominant frames as well as article secondary frames. We define the dominant frame as the most frequent sentence-level frame label in a given document. Secondary frames are all other frames that were assigned to at least one sentence. To reduce noise from near ties, we only assign a dominant frame if the most frequent frame meets both of the following criteria: it appears in at least 3 sentences and it represents at least 40% of all sentences in the document. We make an exception for single-sentence texts (comments), whose frame is automatically considered to be dominant. Articles and comments whose most frequent frame fails to meet these combined thresholds were excluded from further analysis.
@@ 6.3 Topic Extraction
We apply BERTopic (Grootendorst, 2022) separately to SOCC and NYT to extract topics from news articles, assigning each article a probability distribution across all identified topics. [^11] We report the full list of parameters used in Appendix C. We analyzed the initial hierarchical topic structures produced by BERTopic and manually merged closely positioned topics with small inter-topic distances and high thematic similarity. The refined topic model was then applied to reader comments. We ultimately assigned each article and each comment the single most likely topic predicted by BERTopic.
This process was conducted independently for both corpora, to avoid any bias in subsequent analyses. The extracted topics showed substantial thematic overlap across data sets, enabling cross-dataset comparison. We manually identified eleven overarching, semantically cohesive topics which we use in our downstream analysis (Table˜6).
@@
To assess the quality of topic assignments, we randomly sampled 5 articles and 5 comments from each of the 11 topic groups, yielding a total of 110 texts. Two authors of the paper independently reviewed each assigned topic and judged whether it was acceptable or not. We observe a high degree of inter-annotator agreement (96%). The overall accuracy of the model’s predicted topics, calculated against the labels assigned by the annotators was found to be 93% (on both comments and articles).
@@ 6.4 Resulting Data Set
With our framework and validated implementation at hand, we next apply it to analyze framing effects across the SOCC and NYT data sets (Section˜4.1) and eleven selected topics (Section˜6.3). We only include article-comment pairs where both share the same predicted topic (excluding off-topic comments to study framing effects when discussions remain on-topic) and where dominant frames could be extracted for both articles and comments. We also filter out comments shorter than 5 words, and duplicate comments. This final data set consists of nearly 3,000 articles and over 380,000 associated comments, as detailed in Table 6.
@@ 7 Framing Effects across Topics and Outlets
We now address our three research questions (Section 3) asking to what extent dominant article frames stick with readers? How do these trends vary by topic? And, if the dominant frame does not stick, how readers reconstruct the media frame in their comments?
Figure 2: Predicted frame retention rates in reader comments based on regression-derived marginal effects. X- and Y-axes show predicted retention for SOCC and NYT, respectively. Retention varies by frame type, and the predictions correlate strongly across the two media outlets.
@@
7.1 To what extent are dominant article frames retained in reader comments?
In line with our definition of frame retainment as the echoing of the dominant frame of a news article, we first investigate how often the dominant frame kept in its associated direct readers comments. We do so across the New York Times (NYT; USA) and The Globe and Mail (SOCC; Canada).
@@ 7.1 To what extent are dominant article frames retained in reader comments?
Overall, we find that NYT readers demonstrate stronger frame retention (50.7%) compared to SOCC readers (37.3%). Conversely, at least half of the comments (almost two thirds in SOCC) change away from the dominant article frame indicating that readers often do not retain the original primary framing of the article. In section 7.3, we analyze these reframing patterns in depth.
We next investigate how this pattern varies by frame type. We do so through a logistic regression mixed effects model that predicts frame retention (i.e., comment frame being identical to the dominant article frame) as a function of fixed effects for dominant article frame and article topic, and a random effect controlling for article ID, [^12] We control for article ID due to repeated measurements, given that we have multiple comments per article.
$y_{i}\sim\text{topic}_{d}+\text{frame}_{d}+(1|\text{id}_{d}),$ (1)
where the input is a document-response pair $i=(d,r)$, $y_{i}$ is a binary indicator of frame retention ($\text{frame}_{d}{==}\text{frame}_{r}$).
@@
Figure 3: Predicted topic retention rates in reader comments based on regression-derived marginal effects. Points above the diagonal line indicate higher predicted retention in NYT, while points below indicate higher predicted retention in SOCC. Topic names are shortened for readability.
Here, we are interested in the marginal effect of $\text{frame}_{d}$ on $y$ as derived from the fitted model coefficients. Figure 2 displays the frame retention patterns separately for each dominant article frame. Looking at the spread of frames along the x- and y-axis independently, we can see that the extent of retention varies by frame. Specifically, frames such as “Political and Policies”, “Economic”, and “Health and Safety” show high retention rates. In contrast, “Public Opinion”, “Fairness and Equality”, and “Cultural Identity” exhibit substantially lower retention rates. This suggests that readers are more likely to reframe content that is presented through frames that require alignment with personal beliefs or values to be accepted (Chong and Druckman, 2007).
To formally test whether the frame choices in reader comments were influenced by the frames used in the corresponding articles – that is, whether the distribution of frames in comments is not independent of the article’s dominant frame – we conducted chi-squared tests of independence. For each source frame $f$, we tested the null hypothesis that the presence of $f$ in an article is independent of the presence of $f$ in a comment. For all frames, we were able to reject the null hypothesis of independence ($p\ll.001$), although with small to moderate effect sizes (e.g., small effects: “Public Opinion” Cramér’s V = 0.065 in NYT; moderate effects: “Economic” V = 0.357 in SOCC).
Figure 2 additionally reveals that the reframing patterns are remarkably consistent across NYT and SOCC, i.e., all points cluster close to the diagonal line of perfect correlation. This cross-outlet consistency addresses a key limitation of previous single-outlet studies and suggests that our uncovered patterns might reveal a general trend in news reader behavior, rather than being idiosyncratic to a particular outlet or sample.
7.2 How does frame retention vary across topics?
While previous studies have exclusively focused on a single topic and have often produced contradictory results, here we can more rigorously inspect the variation of framing effects across topics under our unified framework.
We explore the effect of topics on framing effects using the same regression framework as in Section 7.1, but this time focusing on the marginal effects of topic on retention from equation (1). This provides predicted frame retention probabilities per topic, controlling for frame type and article-specific effects. The resulting predicted probabilities of frame retention per topic are shown in Figure 3 for NYT (y-axis) and SOCC (x-axis).
We find that predicted framing effects vary considerably by topic, with notable differences between outlets. Several topics show consistent retention patterns across both outlets. Trump & Elections, for instance, demonstrates high predicted retention in both outlets. In contrast, international conflicts like Israel-Palestine and Syria & IS show consistently low retention across both outlets, clustering in the lower portion of the distribution.
Other topics reveal substantial outlet-specific differences. Immigration and Healthcare show notably higher predicted retention in NYT than SOCC, as does coverage of the Russia-Ukraine conflict and LGBT+ Rights, as shown in Figure 3.
These outlet-specific variations, which have not been examined in previous works, confirm that "different kinds of issues are interpreted by the media and by the public in different ways" (Neuman et al., 1992, pg. 17), suggesting that topic-based retention and reframing patterns are more outlet-specific than the frame-based patterns in Figure 2. This finding is in line with the inconsistent results reported in previous small-scale, single-topic studies in the social sciences (Sections 2 and 3).
In this paper, we deliberately remain on the descriptive level and refrain from explaining the observed differences, being cogniscent of many confounding factors that may be at play. However, our framework paves the way for future causal analyses of framing across different topics and news outlets.
7.3 If readers reject the dominant article frame, how do they reframe?
Prior communication studies have used varied approaches to analyze reframing - from binary annotations of the presence of frames (e.g.,Holton et al. (2014); Muñiz et al. (2015)) to dominant frame identification (e.g., Zhou and Moy (2007); Koteyko et al. (2013)). However, none systematically examined how readers reframe when they diverge from the dominant article frame – whether they introduce entirely novel frames or selectively emphasize non-dominant frames already present in the source. Here, we address this question directly.
Comments change away from the dominant article frame 50% (NYT) or 63% (SOCC) of the time, so a natural question is how readers reframe articles. In our analysis, we distinguish between two types of reframing: (a) complete reframing, where the comment frame does not appear in the source article at all, and (b) selective reframing, where comments adopt secondary frames which are present in the source article, but do not dominate the message. In SOCC, 11.7% of comment frames exhibit complete reframing, while 62.8% show selective reframing. In NYT, these fractions are 5.2% and 49.3%. respectively. This shows that, rather than completely accepting or rejecting the dominant source framing, audiences predominantly engage in selective interpretation by adopting secondary frames from the article.
Figure 4 shows reframing patterns for the NYT. Focusing first on complete reframing (Figure˜4(a)), Political framing dominates source articles but commentators introduce diverse alternative frames. In particular readers reframe in terms of value-laden frames including “Morality” (increasing from 0.8% in articles to 26.4% in Comments) or “Cultural Identity”, as well as “Economic” framing. This pattern holds across both NYT and SOCC (as shown in Appendix˜G).
Selective reframing patterns in NYT are shown in Figure 4(b). Here, the frame distributions in the comments align more closely with the dominant article frame distributions, although we again observe an increase in value-laden frames (“Morality”, “Health and Safety”) as well as “Economic” framing.
Topic-specific Reframing
We investigate topic-specific reframing patterns for topics that included at least 100 articles. For instance, in the case of Immigration (Table 7, left), SOCC readers engage in selective reframing by emphasizing both “Cultural Identity” frames when “Political and Policies” dominate the article, and vice versa, suggesting a tight interconnection of these two perspectives. NYT readers show different selective reframing patterns, primarily shifting from “Legality and Crime” frames to “Political and Policies” frames, and from “Political and Policies” to “Economic frames”, indicating a tendency toward more policy-oriented and economic interpretations of immigration coverage. Complete reframing differs substantially (Table 7, right). In NYT, complete reframing shows “Economic” frames being introduced when they were entirely absent from articles. Again, both outlets reframe more objective (economic, political, legal) views into value-laden frames (cultural, morality, fairness). Tables 11 and 12 in the Appendix show selective and complete reframing analysis for additional topics, revealing substantial topic-specific differences.
In conclusion, the results discussed here align with the constructionist view of framing as "interactive, vulnerable and prone to counter-frames" (Gorp, 2007, pg. 76), with audiences engaging in topic-specific patterns of selective retention for frames that resonate with them, whilst actively reframing those that do not.
(a) Complete reframing.
(b) Selective reframing.
Figure 4: Reframing in the New York Times. (a) Complete reframing patterns where readers use frames that never appear anywhere in the original articles. (b) Selective reframing where readers emphasize secondary frames that differ from the dominant article frame but exist elsewhere in the article. Left percentages show dominant article frame distributions; right percentages show comment frame distributions among complete (a) / selective (b) reframing cases.
Selective Reframing Complete Reframing
SOCC Cultural $\rightarrow$ Political Political $\rightarrow$ Cultural
Health $\rightarrow$ Cultural Economic $\rightarrow$ Political
Political $\rightarrow$ Cultural Health $\rightarrow$ Morality
Health $\rightarrow$ Political Economic $\rightarrow$ Cultural
Economic $\rightarrow$ Political Political $\rightarrow$ Fairness
NYT Legality $\rightarrow$ Political Political $\rightarrow$ Economic
Legality $\rightarrow$ Economic Political $\rightarrow$ Legality
Political $\rightarrow$ Economic Legality $\rightarrow$ Cultural
Political $\rightarrow$ Legality Political $\rightarrow$ Cultural
Political $\rightarrow$ Cultural Political $\rightarrow$ Morality